1.The accounting records of Seattle Outlet include the following for January:

Sales Purchases Sales Discounts Freight - In Purchase Returns and Allowances $326,000 260,000 6,000 14,000 2,000

A physical count determined the cost of inventory on hand at January 31 to be $42,000. If gross profit amounts to 25% of net sales, compute the beginning inventory at January 1.

Select one:

a. $10,000

b. $26,000

c. $8,000

d. $24,000

e. $6,000

2.

On 12/31/12, as part of the year-end adjusting journal entries, the Strickland Company accrues three day's wages of $600 ($200 per day). The proper 12/31/12 closing entries are made. No reversing entry is made on 1/1/13. Strickland pays the weekly payroll of $1,000 on 1/2/13. The balance in the Wage Expense account after the 1/2/13 journal entry will be:

Select one:

a. $0

b. $400

c. $600

d. $1,000

e. $1,200

3.

The accounting records of Seattle Outlet include the following for January:

Sales Purchases Sales Discounts Freight - In Purchase Returns and Allowances $326,000 260,000 6,000 14,000 2,000

A physical count determined the cost of inventory on hand at January 31 to be $42,000. If gross profit amounts to 25% of net sales, compute the beginning inventory at January 1.

Select one:

a. $10,000

b. $26,000

c. $8,000

d. $24,000

e. $6,000
Answer #1

1.

Sales = $326,000

Sales discount = $6,000

Net sales = Sales - sales discount

= 326,000-6,000

= $320,000

Gross profit = 25% of net sales

= 320,000 x 25%

= $80,000

Cost of goods sold = Net sales - Gross profit

= 320,000-80,000

= $240,000

Cost of goods sold = Beginning inventory + Purchases - Purchase return + Freight in - Ending inventory

240,000 = Beginning inventory + 260,000-2,000+14,000-42,000

Beginning inventory = $10,000

Correct option is a.

2.

Wages per day = $200

Wages expense for 3 days i.e. $300 was debited on 12/31/12.

Now, on 1/2/13, wages expense will be debited for two days i.e 200 x 2 = $400

The balance in the Wage Expense account after the 1/2/13 journal entry will be: $400

Correct option is (b)

3.

Sales = $326,000

Sales discount = $6,000

Net sales = Sales - sales discount

= 326,000-6,000

= $320,000

Gross profit = 25% of net sales

= 320,000 x 25%

= $80,000

Cost of goods sold = Net sales - Gross profit

= 320,000-80,000

= $240,000

Cost of goods sold = Beginning inventory + Purchases - Purchase return + Freight in - Ending inventory

240,000 = Beginning inventory + 260,000-2,000+14,000-42,000

Beginning inventory = $10,000

Correct option is a.

Kindly comment if you need further assistance.

Thanks‼!